OpenStudy (anonymous):

4x + 2y = 6 x – y = 3 ???

OpenStudy (anonymous):

You need to solve for x and y? Ok, take one equation and then solve that equation for one of the variables. I suggest taking the second one and solving it for x.

OpenStudy (hba):

There are two methods to solve this 1)Elimination. 2)substitution. Which one do you prefer ?

OpenStudy (anonymous):

For this, i think it would be easiest to use substitution: 4x + 2y = 6 x - y = 3 For substitution, you have to isolate one of the variables in one of the equations. the easiest one would be to isolate the x term in the second equation. carry over the y term: x - y = 3 + y + y --------- x = y + 3 Now you can plug this in for x in the first equation 4 ( y + 3) + 2y = 6 Now distribute 4y + 12 + 2y =6 Next, combine like terms, solve like a regular equation 6y + 12 = 6 - 12 - 12 ------------ 6y = -6 y= -1 Now plug this value back in to the first or second equation to get the value of x x - (-6) = 3 x+6 = 3 x= -3